# Macro Graph of Experts for Billion-Scale Multi-Task Recommendation

### v1

@@ 3.2. Macro Graph Bottom
(11) $\widetilde{{\bm{e}}_{v}^{t}}=\operatorname{Lookup}\left(v,\widetilde{{\mathbf{E}}}_{t}\right)$
@@
(13) $\operatorname{MGB}\left(\mathcal{D}_{t},\widetilde{{\mathbf{E}}},\widetilde{\mathcal{N}}\right)=\sum_{{t^{\prime}}\in\mathcal{T}}\operatorname{Gate}_{t^{\prime}}\left({\bm{z}}_{t}\right)\cdot\operatorname{MGE}_{t^{\prime}}\left({\bm{z}}_{t^{\prime}}\right),$
where $\operatorname{Gate}_{t^{\prime}}(\cdot)$ denotes the gating network operation, and $\operatorname{MGE}_{t^{\prime}}(\cdot)$ is defined as Equation 17. $\operatorname{MGB}(\cdot)$ integrates user and item embeddings with their corresponding macro neighbor embeddings, resulting in a more informative representation that reflects the overall macro structure.
@@ 3.3. Macro Graph Expert
(16) $\widetilde{{\mathbf{Z}}}_{v,\widetilde{p}}$ $=\operatorname{softmax}(\frac{{\mathbf{M}}_{\text{sim}}}{\sqrt{d}})\cdot\left({\mathbf{V}}_{v}\cdot\widetilde{{\mathbf{E}}}_{\widetilde{p}}\right),$
@@ 3.4. Macro Prediction Tower
To exploit possible connections between tasks and fully utilize a user’s macro preferences for sparse tasks, we replace the traditional multi-task tower with a macro correlated layer. Specifically, we use a macro cascading readout layer to explore the relationships across different tasks and apply macro task adjustment to align the recommendations with the user’s macro preferences.
@@
Traditional multi-task recommendation models (Ma et al., 2018b; Tang et al., 2020) obtain results in a parallel manner, which ignores the possible connections between different tasks. To address this, we introduce the macro cascading readout module that generates the final output in a cascaded fashion. Specifically, the output of the previous task is fed into the next prediction tower as additional input. Suppose $\hat{y}_{ui}^{(t_{j})}$ is the output of the previous task, then $\hat{y}_{ui}^{(t_{j})}$ can be expressed as:
(21) $\hat{y}_{ui}^{(t_{j})}=\begin{cases}\sigma(\operatorname{MLP}(x^{t_{1}}))\,,&t_{j}=t_{1}\\ \sigma(\operatorname{MLP}(x^{t_{j}})+w_{t_{j-1}}\cdot\hat{y}_{t_{j-1}})\,,&t_{j}\in\mathcal{T}\setminus\{t_{1}\}\\ \end{cases}\,,$
@@
With the goal of recommending more precisely to the user’s individual preferences and enhancing the model’s ability to predict sparse tasks, we propose macro task adjustment to guide the final prediction. First, based on MTMG, we construct a user macro task lookup table according to different task interactions. Specifically, we associate various user’s task interactions with the corresponding macro nodes. For a certain user $u$, the final prediction $\hat{\mathrm{Y}}$ can be expressed as:
(22) $\hat{\mathrm{Y}}={\mathbf{W}}_{u}\cdot{\mathbf{Y}}_{\text{cat}}=\sigma(T_{u})\cdot[\hat{y}_{ui}^{(t_{1})};\;\hat{y}_{ui}^{(t_{2})};\;\ldots]\,,$
where $T_{u}$ is the users’ macro task lookup table, ${\mathbf{W}}_{u}$ represents the macro task adjustment weight corresponding to the user in relation to the target item’s macro node.
@@ 3.5. Model Training
We utilize binary cross-entropy as the loss function during the model’s training. Specifically, with the loss function $\mathcal{L}_{t}$ for a certain task $t$, the objective of MTL is given by:
(23) $\mathcal{L}_{\text{MTL}}=\sum_{t\in\mathcal{T}}\alpha_{t}\mathcal{L}_{t}.$
Then, the overall objective function for MGOE is formulated as:
@@ 4.1.1. Datasets
We conduct extensive experiments on three real-world benchmark datasets: Taobao (Zhu et al., 2018), QB-video (Yuan et al., 2022) and QK-article (Yuan et al., 2022). These datasets include a variety of scenarios, including e-commerce and content-based recommendations, to verify the generality and effectiveness of MGOE in the real-world recommendation. The detailed description and the statistics of these datasets is illustrated in B.1.
@@ 4.1.2. Compared Baselines
To verify the effectiveness of MGOE, we compare with ten relevant representative multi-task models, which can be roughly divided into two categories. (i)Traditional Multi-task Models: MMOE, PLE, ESMM, AITM, STEM, MoME. (ii) Graph-based Methods: MoGENet, MMoCEG, POGCN, MacGNN. We leave the details of these baseline models in B.2.
@@ 4.1.3. Implementation Details
All baselines and MGOE is implemented based on Pytorch. In the offline experiments, the embedding size fix to 10 and use Xavier (Glorot and Bengio, 2010) for initialization for all models. For MGOE, we set the batch size to 1024 and use Adam (Kingma, 2014) for optimization. The learning rate is searched form {1e-2, 5e-2, 1e-3, 5e-3} and the regularization term is searched form {1e-4, 5e-5, 1e-5}.
@@ 4.1.4. Evaluation Metrics
We adopt two widely used multi-task evaluation metrics, AUC (Fawcett, 2006) and Logloss (Chen et al., 2016), for a comprehensive evaluation of the models. Higher values of AUC indicate better prediction performance of the model, while the lower values of Logloss imply better prediction accuracy. Higher values of AUC indicate better CTR prediction performance of the model, while the lower values of Logloss indicate better CTR prediction performance. We run the experiments five times with different random seeds and take the average result to avoid extreme cases.
@@ 4.2. Main Results (RQ1)
Introducing graph information can effectively improve the performance of the models. Based on the results, we can observe that graph-based models, like MMoCEG and MacGNN, generally yield better results compared to traditional multi-task models. This implies that traditional multi-task models treat users and items in isolation, overlooking the relationship between them. Conversely, integrating the information of the user-item bipartite graph can mirror the more profound interactions between users and items. This is of great significance for enhancing the overall recommendation performance.
Introducing graph information from the micro level can not fully utilize the task patterns. By comparing the micro-graph based methods and macro-graph models in Table 1, it can be observed that macro-graph models are capable of achieving relatively superior performance in multi-task recommendation compared to micro-graph ones. This indicates that the traditional microscopic graph fails to effectively translate the diverse tasks of users towards items into enhanced model performance. Consequently, there is a need to integrate these user task patterns in a novel manner and consider these user and item nodes from a macroscopic perspective, which serves as the impetus for proposing MTMG.
@@
Taobao MGOE 0.6981 0.2042
w/o preference 0.6912 0.2054
w/o adjustment 0.6854 0.2059
w/o cascading 0.6907 0.2054
QB-video MGOE 0.7839 0.1123
w/o preference 0.7806 0.1130
w/o adjustment 0.7659 0.1211
w/o cascading 0.7716 0.1164
@@ 4.3. Ablation Study (RQ2)
  • The macro prediction tower facilitates the macro information into more accurate recommendations. Removing the macro cascading readout module leads to a decline in performance. This indicates the necessity of representing the transfer of macro knowledge across different tasks. When the macro task adjustment is removed, the performance deteriorates, as the macro nodes preserve a wealth of users’ preference information.
Figure 3. Parameter study of $\gamma$ and $\widetilde{n}$ on Taobao and QB-video.
@@ 4.4. Hyper-parameter Study (RQ3)
We investigate three key hyperparameters in MGOE: (1) the temperature parameter $\gamma$ for weighting macro nodes; (2) the number of macro nodes in MTMG, denoted as $\widetilde{n}$; (3) the weight $w_{v}$ in preference-based grouping. To explore the impact of these hyperparameters on MGOE, we conduct an analysis by varying the settings of the key hyperparameters on the Taobao and QB-video datasets. While examining $\gamma$, $\widetilde{n}$, and $w_{v}$, we keep all other hyperparameters fixed. The results are shown in Figure 3 and Figure 4.
@@
Effect of Macro Node Number. We vary the number of macro nodes from 5 to 30 in steps of 5. From the results presented in the bottom half of Figure 3, we can observe that an overly small number of macro nodes tends to lead to overly coarse segmentation, resulting in poor outcomes. Additionally, selecting a relatively appropriate number of macro nodes can yield satisfactory performance. For both Taobao and QB-video, when the number of macro nodes reaches 20, MGOE achieves the best performance.
Effect of Preference-Based Grouping Weight. We set the value of $w_{\text{high}}$ to either 2 or 1, and $w_{\text{low}}$ to either 1 or 0.5. Thus, we can represent the four combinations as follows: $w_{1}$=(2, 0.5), $w_{2}$=(2, 1), $w_{3}$=(1, 0.5) and $w_{4}$=(1, 1), where the first value corresponds to $w_{\text{high}}$ and the second to $w_{\text{low}}$. From the results presented in Figure 4, we can observe that in both datasets, assigning larger weights to active micro nodes and relatively smaller weight to non-active micro nodes generally leads to better performance.
@@ 4.5. Efficiency Study (RQ4)
Since multi-task recommendation models are often deployed in real-time scenarios where multiple tasks need to be handled simultaneously, the computational efficiency of these models is a crucial evaluation factor (Wang et al., 2019b). Hence, to verify the efficiency of our proposed MGOE, we compare the average inference time between MGOE and several representative and graph-based multi-task baselines: expert-based model MMoE and task-specific embedding scheme STEM, graph-based method POGCN, multi-channels paradigm MoGENet and graph MOE-based MMoCEG.
The comparison results are presented in Figure 5. From the figure, we have the following observations: (i) Our proposed MGOE is nearly as efficient as the simplest MMoE model and on par with STEM in terms of efficiency. While maintaining the same inference time as these models, our model incorporates high-order graph information, which significantly enhances the performance of multi-task recommendation. (ii) Our model has a much shorter inference time compared with traditional graph-based multi-task models. MoGENet requires almost four times the inference time of our model, indicating that our approach to integrating high-order graph information is far more efficient, enabling faster and more responsive multi-task recommendations, which is of great significance for real-time applications in large-scale recommendation systems.
@@ 5. Conclusion
Appendix A RELATED WORK
A.1. Multi-task Learning
Multi-task learning aims to enhance the model’s generalization ability by training multiple tasks concurrently within the same model. The core of this method lies in the knowledge that can be shared among different tasks, enabling them to benefit from one another. In the early days of deep learning, the hard sharing mechanism of parameters was commonly employed. It could be applied to all hidden layers for all tasks, while maintaining task-specific output layers. The parameter soft sharing mechanism, on the other hand, allows each task to have its own model and parameters, thereby achieving better performance. The cross-stitch network utilizes the cross-stitch unit to take a task-specific network and linearly combine it with the previous output by leveraging knowledge learned from other tasks. However, in recommendation systems, parameter sharing between tasks that have poor correlation can lead to a seesaw phenomenon. To solve the issues of data sparsity (DS) and sample selection bias (SSB), ESSM (Ma et al., 2018a) simultaneously optimizes two related tasks in a sequential manner. MOE (Jacobs et al., 1991) employs a gated network to transform the original parameters shared by all samples into multiple sets of parameters, where each set is called an expert. MMOE (Ma et al., 2018b) proposes a multi-expert network to address the multi-task learning problem, with the hope that the model could automatically learn to assign expert parameters based on the relationship of the underlying task. PLE (Tang et al., 2020) divides the model parameters into a private part and a shared part, which improves the robustness of multi-task learning and mitigates the negative impact between private knowledge and shared knowledge.
A.2. Graph Learning for Recommendation
In recent years, an increasing number of studies have attempted to integrate graphs into recommendation systems in order to enhance performance. NGCF (Wang et al., 2019a) typically incorporates higher-order graph information into traditional collaborative filtering. To achieve superior performance, LightGCN (He et al., 2020) removes non-linear operations and feature transformations, which are found to be redundant through experimentation. These methods have been widely adopted in industrial item recall scenarios. However, due to their long inference times, these methods are difficult to apply directly to multi-task learning. Subsequently, some approaches attempt to incorporate graph information into the CTR prediction scenario. MoGENet (Hu et al., 2022) proposes a novel multi-channel graph neural network to jointly model high-order information using the user-item bipartite graph. MMoCEG (Zhang et al., 2023b) utilizes a graph-based structure to share information across regions and employs contrastive learning to prevent expert representation degeneration. GLSM (Sun et al., 2022) employs an adaptive fusion mechanism to fuse information from long-term and short-term behaviors. GMT (Min et al., 2022) effectively integrates different types of interactions between local neighborhood nodes from both explicit and implicit perspectives to generate highly representative embeddings for the target user-item pair. MacGNN (Chen et al., 2024) attempts to cluster similar micro nodes into a macro node to reduce computational complexity.
A.3. Click-Through Rate Prediction
Click-Through Rate (CTR) Prediction is one of the core algorithms in the recommendation system  (Zhu et al., 2021; Zhang et al., 2021), which predicts whether a user will click on an advertisement or not. The CTR prediction is influenced by the user’s historical click-through rate and behavior, and it is trained on vast amounts of historical data by combining various factors and characteristics. Traditional shallow CTR models, such as Logistic Regression (LR), Degree-2 Polynomial Mappings (Poly2), and Factorization Machine (FM) (Rendle, 2010), were widely used in the early days due to their lightweight nature and strong interpretability. With the development of deep learning, modern deep learning methods have achieved significant improvements. FNN uses FM in the embedding layer to generate hidden vectors as initialization values for embeddings. PNN (Qu et al., 2016) adds a product layer between the embedding layer and the MLP. By adding an attention mechanism between the embedding layer and the MLP, DIN (Zhou et al., 2018) can learn from the user’s past behavior. DIEN (Zhou et al., 2019) incorporates the ARGRU sequence model to model the evolution process of interest. DeepFM (Guo et al., 2017) replaces LR in the traditional wide part with FM, and DCN (Wang et al., 2017) further replaces FM with a cross layer network to handle feature crossover. DCNv2 (Wang et al., 2021) uses low-rank mixing to strike a balance between performance and delay, thereby achieving state-of-the-art (SOTA) results.
Appendix B EXPERIMENTAL DETAILS
B.1. Dataset Details
  • Taobao (Zhu et al., 2018) is a user behavior dataset provided by Alibaba. Taobao is one of the most famous online shopping malls in China, the dataset includes the interaction behaviors between 28,910 users and 65,087 items from November to December 2017, including clicking, favoring, adding to cart and buying. The Taobao dataset is available on the Alibaba Tianchi page [^1] https://tianchi.aliyun.com/dataset/649.
  • QB-video (Yuan et al., 2022) is a large-scale real-world dataset obtained from the recommendation data of two content platform applications of Tencent. It has over 5 million users and 140 million interactions. We select a subset of video scenes, which mainly consists of 28,803 users and 15,540 items, including four basic behaviors: clicking, liking, sharing and following. The QB-video dataset can be obtained from the information page of the Tencent dataset [^2] https://github.com/yuangh-x/2022-NIPS-Tenrec/tree/master.
  • QK-article (Yuan et al., 2022) is a dataset in Tenrec under the article scene provided by Tencent. It records the behaviors of 52,052 users and 30,273 products, which mainly fall into four categories: reading, liking, favoring and following. The dataset can be accessed on the repository [^3] https://github.com/yuangh-x/2022-NIPS-Tenrec/tree/master.

### v4

@@ 3.2. Macro Graph Bottom
(12) $\widetilde{{\bm{e}}_{v}^{t}}=\operatorname{Lookup}\left(v,\widetilde{{\mathbf{E}}}_{t}\right),$
@@
(14) $\operatorname{MGB}\left(\mathcal{D}_{t},\widetilde{{\mathbf{E}}},\widetilde{\mathcal{N}}\right)=\sum_{{t^{\prime}}\in\mathcal{T}}\operatorname{G}_{t^{\prime}}^{(t)}\left(\operatorname{MGE}_{t}({\bm{z}}_{t})\right)\cdot\operatorname{MGE}_{t^{\prime}}\left({\bm{z}}_{t^{\prime}}\right),$
where $\operatorname{G}_{t^{\prime}}^{(t)}(\cdot)$ denotes the gating network operation of task $t$, and $\operatorname{MGE}_{t^{\prime}}(\cdot)$ is defined as Equation 18. $\operatorname{MGB}$ integrates user and item embeddings with their corresponding macro neighbor embeddings, resulting in a more informative representation that reflects the overall macro structure.
@@ 3.3. Macro Graph Expert
[…]
@@ 3.4. Macro Prediction Tower
To exploit possible connections between tasks and fully utilize a user’s macro preferences for sparse tasks, we replace the traditional multi-task tower with a macro correlated layer. Specifically, we use a macro cascading readout module to explore the relationships across different tasks and apply macro task adjustment to align the recommendations with the user’s macro preferences.
@@
Traditional multi-task recommendation models (Ma et al., 2018a; Tang et al., 2020) obtain results in a parallel manner, which ignores the possible connections between different tasks. To address this, we introduce the macro cascading readout module that generates the final output in a cascaded fashion. Specifically, the output of the previous task is fed into the next prediction tower as additional input. Suppose $\hat{y}_{ui}^{(t_{j-1})}$ is the output of the previous task, then $\hat{y}_{ui}^{(t_{j})}$ can be expressed as:
(22) $\hat{y}_{ui}^{(t_{j})}=\begin{cases}\sigma(\operatorname{MLP}(x^{t_{1}}))\,,&t_{j}=t_{1}\\ \sigma(\operatorname{MLP}(x^{t_{j}})+w_{t_{j-1}}\cdot\hat{y}_{ui}^{(t_{j-1})})\,,&t_{j}\in\mathcal{T}\setminus\{t_{1}\}\\ \end{cases}\,$,
@@
With the goal of recommending more precisely to the user’s individual preferences and enhancing the model’s ability to predict sparse tasks, we propose macro task adjustment to guide the final prediction. First, based on MTMG, we construct a user macro task lookup table according to different task interactions. Specifically, we associate various user’s task interactions with the corresponding macro nodes. For a certain user $u$, the final prediction $\hat{\mathrm{Y}}_{ui}$ can be expressed as:
(23) $\hat{\mathrm{Y}}_{ui}={\mathbf{W}}_{u}\odot{\mathbf{Y}}_{\text{cat}}=\sigma(T_{u})\odot[\hat{y}_{ui}^{(t_{1})};\;\hat{y}_{ui}^{(t_{2})};\;\ldots]\,$,
where $T_{u}$ is the users’ macro task lookup table, ${\mathbf{W}}_{u}$ denotes the user-specific macro task adjustment weights. The prediction is computed via the Hadamard product (element-wise multiplication) $\odot$.
@@ 3.5. Model Training
We utilize binary cross-entropy as the loss function during the model’s training. Specifically, with the loss function $\mathcal{L}_{t}$ for a certain task $t$, the objective of MTL $\mathcal{L}_{\text{MTL}}$ is given by Equation 2. Then, the overall objective function for MGOE is formulated as:
@@ 4.1.1. Datasets
We conduct extensive experiments on three real-world benchmark datasets: Taobao (Zhu et al., 2018), QB-video (Yuan et al., 2022) and QK-article (Yuan et al., 2022). The detailed description and statistics of these datasets are illustrated in Appendix A.1.
@@ 4.1.2. Compared Baselines
To verify the effectiveness of MGOE, we compare with thirteen relevant representative models, which can be roughly divided into three categories. (i) Traditional Multi-task Models: MMoE, PLE, ESMM, AITM, STEM, MoME. (ii) Graph-based Methods: MoGENet, MMoCEG, POGCN, MacGNN. (iii) Single-Task Learning Methods: DeepFM, DIN, DIEN. We leave the details of these baseline models in Appendix A.2.
@@ 4.1.3. Implementation Details
All baselines and MGOE are implemented based on Pytorch. In the offline experiments, the embedding size fix to 10 and use Xavier (Glorot and Bengio, 2010) for initialization for all models. For MGOE, we set the batch size to 1024 and use Adam (Kingma, 2014) for optimization. The learning rate is searched from {1e-2, 5e-2, 1e-3, 5e-3} and the regularization term is searched from {1e-4, 5e-5, 1e-5}.
@@ 4.1.4. Evaluation Metrics
We adopt two widely used multi-task evaluation metrics, AUC (Fawcett, 2006) and Logloss (Chen et al., 2016), for a comprehensive evaluation. Higher values of AUC indicate better prediction performance of the model, while the lower values of Logloss imply better prediction accuracy. We run the experiments five times with different seeds with standard deviation to prevent extreme cases, the result is shown in Table 6 in Appendix A.3. Additionally, the results of the performance of single-task learning (STL) models are included in Table 7 in Appendix A.4.
@@ 4.2. Main Results (RQ1)
Introducing graph information can effectively improve the performance of the models. Based on the results, we can observe that graph-based models, like MMoCEG and MacGNN, generally yield better results compared to traditional multi-task models. This implies that traditional multi-task models treat users and items in isolation, overlooking the relationship between them. Conversely, integrating information from the user-item bipartite graph can mirror the more profound interactions between users and items. This is of great significance for enhancing the overall recommendation performance.
Introducing graph information from the micro level can not fully utilize the task patterns. By comparing the micro graph based methods and macro graph models in Table 1, it can be observed that macro graph models are capable of achieving relatively superior performance in multi-task recommendation compared to micro graph ones. This indicates that the traditional microscopic graph fails to effectively translate the diverse tasks of users towards items into enhanced model performance. Consequently, there is a need to integrate these user task patterns in a novel manner and consider these user and item nodes from a macroscopic perspective, which serves as the impetus for proposing MTMG.
@@
Taobao MGOE 0.6936±0.0040 0.2057±0.0010
w/o preference 0.6919±0.0048 0.2060±0.0009
w/o adjustment 0.6836±0.0047 0.2080±0.0031
w/o cascading 0.6927±0.0058 0.2066±0.0009
QB-video MGOE 0.7715±0.0114 0.1124±0.0020
w/o preference 0.7700±0.0201 0.1127±0.0023
w/o adjustment 0.7513±0.0162 0.1140±0.0040
w/o cascading 0.7640±0.0135 0.1142±0.0033
@@ 4.3. Ablation Study (RQ2)
  • The macro prediction tower facilitates the macro information into more accurate recommendations. When the macro task adjustment is removed, the performance deteriorates, as the macro nodes preserve a wealth of users’ preference information. Removing the macro cascading readout module leads to a decline in performance. This indicates the necessity of representing the transfer of macro knowledge across different tasks.
@@
Figure 4. Parameter study of $\gamma$ and $\widetilde{m}$ on Taobao and QB-video.
@@ 4.4. Hyper-parameter Study (RQ3)
We investigate three key hyperparameters in MGOE: (1) the weight $w_{v}$ in preference-based grouping; (2) the temperature parameter $\gamma$ for weighting macro nodes; (3) the number of macro user nodes in MTMG, denoted as $\widetilde{m}$. We conduct grid search within a small range over key hyperparameters on the Taobao and QB-video datasets. While examining $w_{v},\gamma,\widetilde{m}$, we keep all other hyperparameters fixed. The results are shown in Figure  3 and Figure 4.
Effect of Preference-Based Grouping Weight. We set the value of $w^{+}$ to either 2 or 1, and $w^{-}$ to either 1 or 0.5. Thus, we can represent the four combinations as follows: $w_{1}$=(2, 0.5), $w_{2}$=(2, 1), $w_{3}$=(1, 0.5) and $w_{4}$=(1, 1), where the first value corresponds to $w^{+}$ and the second to $w^{-}$. From the results presented in Figure 3, we can observe that in both datasets, assigning larger weights to active micro nodes and relatively smaller weights to inactive micro nodes generally leads to better performance.
@@
Effect of Macro Node Number. We vary the number of macro user nodes from 5 to 30 in steps of 5. From the results presented in the bottom half of Figure 4, we can observe that an overly small number of macro user nodes tends to lead to overly coarse segmentation, resulting in poor outcomes. Additionally, selecting a relatively appropriate number of macro nodes can yield satisfactory performance. For both Taobao and QB-video, when the number of macro user nodes reaches 20, MGOE achieves the best performance.
@@ 4.5. Efficiency Study (RQ4)
Since multi-task recommendation models are typically deployed in real-time scenarios, computational efficiency is a critical metric (Wang et al., 2019a). To verify the efficiency, we compare the average inference time, model parameter size, GPU memory usage, and training time between MGOE and several representative multi-task baselines.
4.5.1. Average Inference Time
The comparison results are presented in Figure 5. From the figure, we have the following observations: (i) MGOE is nearly as efficient as the simplest expert-based model and on par with STEM in terms of efficiency. While maintaining the same inference time as these models, our model incorporates high-order graph information, which significantly enhances the performance of multi-task recommendation. (ii) MGOE has a much shorter inference time compared to graph-based models. Specifically, on the Taobao dataset, MGOE’s inference time is 55.83% faster than POGCN, 72.64% faster than MoGENet and 45.13% faster than MMoCEG. On the QK-article dataset, MGOE achieves an improvement of 56.31%, 80.49%, and 35.58% respectively. This shows that our approach to integrating high-order graph information is much more computationally efficient, enabling faster and more responsive recommendations. (iii) Among all the models, those without expert networks (MacGNN, ESMM) have the fastest inference speed, as the main source of inference time comes from the expert architecture.
Table 3. Model parameter size, GPU memory usage and training time comparison on Taobao dataset.
Models Parameter Size/M GPU Usage/MiB Training Time/s
MGOE 5.04 1194 258
MoGENet 11.57 2992 418
MMoCEG 12.22 1486 391
MMoE 3.21 1056 333
PLE 13.29 1302 957
STEM 13.59 1480 438
MacGNN 1.50 1080 320
4.5.2. Model Parameter Size
In Table 3, we provide a detailed comparison of model parameter size. Among expert-based models, MGOE demonstrates advantages with a parameter size of 5.04M. Compared to other models such as PLE (13.29M) and MMoCEG (12.22M), MGOE has a significantly reduced parameter size, specifically reduced by approximately 62.08% and 58.76% respectively. Although some models like MacGNN have a smaller parameter size (1.50M), MGOE maintain a balance between model complexity and performance, making it more suitable for real-world applications.
4.5.3. GPU Memory Usage and Training Time
Table 3 presents the GPU memory usage (in MiB) and the training time (in seconds) of MGOE and other competing models on the Taobao dataset. Specifically, MGOE requires only 1194 MiB of GPU memory, which represents a reduction of approximately 60.10% compared to MoGENet and 19.66% compared to MMoCEG. Additionally, our model’s GPU usage is also lower than that of PLE and STEM, and slightly higher than that of MMoE and MacGNN. These results demonstrate the efficiency of our model in terms of computational resource usage. Besides, the training time of MGOE is remarkably shorter compared to other methods, indicating that MGOE can converge smoothly.
@@ 4.6. Online Evaluation (RQ5)
5. Related Works
5.1. Multi-Task Learning
Multi-task learning aims to enhance the model’s generalization ability by training multiple tasks concurrently within the same model. In the early days of deep learning, the hard sharing mechanism of parameters was commonly employed. It could be applied to all hidden layers for all tasks, while maintaining task-specific output layers. The parameter soft sharing mechanism, on the other hand, allows each task to have its own model and parameters, thereby achieving better performance. To solve the issues of data sparsity (DS) and sample selection bias (SSB), ESSM (Ma et al., 2018b) simultaneously optimizes two related tasks in a sequential manner. MOE (Jacobs et al., 1991) employs a gated network to transform the original parameters shared by all samples into multiple sets of parameters, where each set is called an expert. MMOE (Ma et al., 2018a) proposes a multi-gate mixture-of-experts framework to explicitly learn to model task relationships. PLE (Tang et al., 2020) divides the model parameters into a private part and a shared part, which improves the robustness of multi-task learning and mitigates the negative transfer problem.
5.2. Graph Learning for Recommendation
In recent years, graphs have been increasingly incorporated into recommendation systems to enhance performance (Ying et al., 2018; Bei et al., 2025; Zhang et al., 2025; Huang et al., 2025). NGCF (Wang et al., 2019b) introduces higher-order connectivity into collaborative filtering, while LightGCN (He et al., 2020) further improves efficiency by removing non-linear transformations. However, the long inference time of such models limits their direct applicability to multi-task learning, motivating subsequent efforts to integrate graph structures into CTR prediction. MoGENet (Hu et al., 2022) adopts a multi-channel GNN to jointly model high-order signals on the user–item bipartite graph. MMoCEG (Zhang et al., 2023a) leverages a graph-based expert-sharing framework with contrastive learning to alleviate representation degeneration. GLSM (Sun et al., 2022) employs adaptive fusion to combine long-term and short-term behavioral information. GMT (Min et al., 2022) models heterogeneous local interactions to generate expressive embeddings for target user–item pairs. MacGNN (Chen et al., 2024) reduces computational cost by clustering similar micro-nodes into macro-nodes.
@@ 6. Conclusion
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (Grant Nos. 62272200, 62502008, and 7250129), in part by the Funding Scheme for Research and Innovation of FDCT (Grant No. 0021/2025/ITP1), and in part by the Youth Independent Innovation Science Fund of the National University of Defense Technology (Grant No. ZK25-61).
Appendix A Experimental Details